# A theory of ecological invasions and its implications for eco-evolutionary dynamics

### v1

@@ V Validation using experimental data
To test whether our framework could successfully predict invasion outcomes in experimental communities, we analyzed three datasets spanning diverse ecological systems (results in Fig. 5). Each dataset included species abundance measurements collected under consistent environmental conditions, with different combinations of species seeded in the ecosystem. Each ecosystem contained 4–5 species spanning different length scales and trophic levels: from plants to herbivore-algae to ciliated protists.
@@
Even though these were not originally designed as invasion experiments, the large number of species combinations allowed us to probe invasion outcomes using a comparative approach. We took pairs of species combinations that differed from each other by the introduction of exactly one species. We then applied our framework viewing this one species as an invader. Thus, the community where this species was not introduced was the pre-invasion community, and the one where it was introduced was the post-invasion community. The large number of species combinations and relatively low community diversity allows one to infer species interactions matrices (assuming GLV dynamics) for each ecosystem. Thus, these experiments provide all the information needed to implement our framework, and comprise an ideal to test of our theory.
@@
Applying our theoretical framework to all three datasets, we predicted: (a) whether the invader would successfully invade; if yes, the post-invasion (b) invader abundance and (c) resulting shift in community abundances (for details, see Methods in Appendix I). Fig. 5 shows results for all three datasets: (a) plant communities from the southeastern United States (data from Kuebbing et al. [38]); (b) herbivore-algae communities (data from Rakowski and Cardinale [39]); and (c) ciliated protist communities grown across a range of temperatures (15°C to 25°C; data from Pennekamp et al [40]). We found that predictions broadly agreed with measurements across all datasets. Specifically, while the median of our predictions agreed with the median of measurements, the size of the error bars varied across datasets. Generally, prediction errors decreased with the number of replicates. Errors were smallest for the plant community dataset (Fig. 5a), which had the largest number of replicates (10); and lowest for herbivore-algae communities (Fig. 5b) with the lowest number of replicates (5) . Further, deviations from measurements were more likely for species with low abundances (e.g., see points on the bottom-right of both panels of Fig. 5c). These deviations stem from several sources of error: measurement errors, biological variability, uncertainty in fitted interaction strengths, as well as GLV dynamics being an approximation of the true dynamics. Despite these assumptions, our framework demonstrates that invasion outcomes can be predicted and understood in diverse ecosystems spanning a range of environments, size scales, and trophic levels.
@@ VI Discussion
A fundamental insight from our theory is the identification of a “dressed invasion fitness” as the central quantity governing invasion outcomes. Unlike conventional “naive” invasion fitness [34, 41] which only measures the initial growth rate of an invader in the resident community, the dressed invasion fitness incorporates ecological feedbacks between the invader and the resident community. We show that this distinction is crucial for accurately predicting both invasion success and the resulting community structure. This is because mathematically, the effect of an invader abundance decomposes into a contribution proportional to its dressed invasion fitness and a contribution from invasion-induced extinctions. Both these contributions incorporate ecological feedbacks between the invader and the full community into which it invades.
A direct implication of our framework is that eco-evolutionary dynamics in complex communities follow a pattern where successful mutants predominantly replace low-abundance community members rather than their parent species [10]. This stands in contrast to the canonical view from population genetics — where displacement occurs among closely related lineages — as well as pairwise niche overlap metrics [14]. The mathematical basis for this phenomenon lies in our self-consistency equations, particularly in how the dressed invasion fitness incorporates global feedbacks mediated by the entire community structure. The resulting extinction probabilities decay exponentially with pre-invasion abundance, effectively shielding abundant species while rendering rare species vulnerable regardless of their relatedness to the invader. This observation highlights how ecological feedbacks fundamentally reshape evolutionary trajectories in ways that cannot be captured by traditional models that ignore extinction-induced feedback.
@@ Acknowledgments
Appendix A Derivation of linear response theory for Lotka-Volterra model
@@ A.1 Primer: Environmental perturbations without extinctions
$\vec{X}^{\text{new}}_{\mathcal{S}}$ $=A_{\mathcal{S}\mathcal{S}}^{-1}(\vec{r_{\mathcal{S}}}-A_{\mathcal{S}0}{X}_{0})$ (18)
@@
$\vec{r_{\mathcal{S}}}-A_{\mathcal{S}\mathcal{S}}\vec{X}^{\text{old}}_{\mathcal{S}}-A_{\mathcal{S}\mathcal{E}}\vec{X}_{\mathcal{E}}$ $=0$ (19)
$\vec{X}^{\text{old}}_{\mathcal{S}}$ $=A_{\mathcal{S}\mathcal{S}}^{-1}(\vec{r_{\mathcal{S}}}-A_{\mathcal{S}\mathcal{E}}\vec{X}_{\mathcal{E}})$ (20)
@@
$\delta\vec{X}_{\mathcal{S}}$ $=\vec{X}^{\text{new}}_{\mathcal{S}}-\vec{X}^{\text{old}}_{\mathcal{S}}$ (21)
$=A_{\mathcal{S}\mathcal{S}}^{-1}(-A_{\mathcal{S}0}{X}_{0}+A_{\mathcal{S}\mathcal{E}}\vec{X}_{\mathcal{E}})$ (22)
@@ A.4 Extended framework for multiple invaders and environmental perturbations
$\vec{r}^{\text{inv}}=-A_{\mathcal{S}\mathcal{I}}\vec{X}_{\mathcal{I}},\quad\vec{r}^{\text{ext}}=A_{\mathcal{S}\mathcal{E}}\vec{X}_{\mathcal{E}}$
@@
Substituting (29) into (LABEL:eq:invaderx) like before, we obtain the generalized version of Eq. (5)
Figure S1: Schematized equations showing how our theory partitions an ecosystem into invaders, survivors and extinct species. (a) shows how the structure of interactions changes in the Lotka-Volterra model before and after an invasion, and (b) shows how the community shifts and invader abundance is related to each part of the system shown in Eq. (27)
@@
Appendix B General linear response theory
@@
where $X_{i}$ represents the abundance of each species, and $\vec{r}$ parameterizes the environmental influences on growth. The function $f_{i}(X_{i})$ governs species extinction dynamics. It is positive for surviving species, and it captures the concept of extinction for biotic species as $f_{i}(X_{i})$ must satisfy one of the following conditions:
$f_{i}(0)$ $=0,\quad\text{for biotic species},$ (37)
@@
where the second condition corresponds to the stability to re-introduction by extinct species. This tells us that instead of linearizing the complete dynamics in Eq.(36), it is more natural to linearize the steady-state condition if we consider perturbations to the ecosystem.
We first study the scenario where all species survive, following the first condition in Eq.(38). We want to study the abundance change to each species, $\vec{X}\to\vec{X}^{*}+\delta\vec{X}$, in response to perturbation of environmental parameter $\vec{r}\to\vec{r}+\delta\vec{r}$. The new steady state needs to satisfy
@@
Define the susceptibility matrices of the steady state condition,
$\mathcal{A}_{ij}=\left.\frac{\partial g_{i}(\vec{X},\vec{r})}{\partial X_{j}}\right|_{\vec{X}=\vec{X}^{*}},\quad\mathcal{B}_{ij}=\left.\frac{\partial g_{i}(\vec{X},\vec{r})}{\partial r_{j}}\right|_{\vec{X}=\vec{X}^{*}}$ (40)
Taylor expanding the new steady state condition in Eq.(39) to first order results in
$\mathcal{A}\delta\vec{X}+\mathcal{B}\delta\vec{r}=0$ (41)
Further simplifying by labeling the matrix product $A^{-1}=-\mathcal{A}^{-1}\mathcal{B}$. The abundance shift should satisfy
$\delta\vec{X}={A}^{-1}\delta\vec{r}$ (42)
To summarize, to predict the linear response of the steady state in any nonlinear model, the first step is to compute the susceptibility matrix, and compute the matrix inverse given by $A^{-1}$. Once this matrix is determined, the problem simplifies and becomes equivalent to a linear model with
$g_{i}(\vec{X},\vec{r})=r_{i}-\sum_{j}A_{ij}X_{j},$ (43)
which has the exact same form as the steady state condition for the general Lotka-Volterra model. This simple exercise of Taylor expansion justifies the use of the Lotka-Volterra model despite its simplicity - it captures the linear response of ecosystems near their steady states. Finally, the result maps to the one derived for the general Lotka-Volterra model in Eq.(6). A similar derivation for Eq.(5) will require solving only the equation for invader steady state abundance, i.e., $g_{i\in\mathcal{I}}(\vec{X}^{\text{new}},\vec{r}^{\text{new}})=0$.
Appendix C Extinct species have negative invasion fitness
@@
$\frac{d}{dt}(x_{i}+\delta x_{i})=(x_{i}+\delta x_{i})(k_{i}-\sum_{j\neq i}A_{ij}x_{j}-A_{ii}\delta x_{i})$
@@
$\frac{d\delta x_{i}}{dt}=\delta x_{i}(k_{i}-\sum_{j\neq i}A_{ij}x_{j}-A_{ii}\delta x_{i})$
@@
$\frac{d\delta X_{i}}{dt}=g_{i}(\vec{X},\vec{r})\frac{\partial f_{i}}{\partial X_{i}}\delta X_{i}+f_{i}(X_{i})\frac{\partial g_{i}}{\partial X_{i}}\delta X_{i}$
@@
$\left.g_{i}(\vec{X},\vec{r})\frac{\partial f_{i}}{\partial X_{i}}\right|_{X_{i}=0}<0+\left.f_{i}(X_{i})\frac{\partial g_{i}}{\partial X_{i}}\right|_{X_{i}=0}<0$
@@
Appendix D Derivation of linear response in CRMs

### v2

@@ V Validation using experimental data
To test whether our framework could successfully predict invasion outcomes in stable experimental communities, we analyzed six datasets spanning diverse ecological systems (three shown in Fig. 5; remaining in Fig. S7). Each dataset included species abundance measurements collected under consistent environmental conditions, with different combinations of species seeded in the ecosystem. Each ecosystem contained 4–5 species spanning different length scales and trophic levels: from plants to herbivore-algae to ciliated protists.
@@
Even though these were not originally designed as invasion experiments, the large number of species combinations allowed us to probe invasion outcomes using a comparative approach. We took pairs of species combinations that differed from each other by the introduction of exactly one species. We then applied our framework viewing this one species as an invader. Thus, the community where this species was not introduced was the pre-invasion community, and the one where it was introduced was the post-invasion community. The large number of species combinations and relatively low community diversity allows one to infer species interactions matrices for each ecosystem, done separately by fitting a GLV model to the data [42, 43]. In some datasets, the presence of experimental replicates allow us to estimate error bars. These experiments provide the information needed to implement our framework, and comprise an appropriate test of our theory.
@@
Applying our theoretical framework to all datasets, we predicted: (a) whether the invader would successfully invade; if yes, the post-invasion (b) invader abundance and (c) resulting shift in community abundances (for details, see Methods in Appendix I). Fig. 5 shows results for three datasets: (a) plant communities from the southeastern United States (data from Kuebbing et al. [39]); (b) herbivore-algae communities (data from Rakowski and Cardinale [40]); and (c) ciliated protist communities grown across a range of temperatures (15°C to 25°C; data from Pennekamp et al [41]). Fig. S7 shows three additional plant datasets from a series of distinct biodiversity-ecosystem functioning (BEF) experiments [44, 45, 46]. We found that predictions broadly agreed with measurements across all datasets. Specifically, while the median of our predictions agreed with the median of measurements, the size of the error bars varied across datasets. Generally, prediction errors decreased with the number of replicates, suggesting that experimental noise in the datasets was the dominant source of error. Errors were smallest for the plant community dataset (Fig. 5a), which had the largest number of replicates (10); and highest for herbivore-algae communities (Fig. 5b) with the lowest number of replicates (5). The additional datasets (Fig. S7) had no replicates, so we could not estimate error bars for them. Further, deviations from measurements were more likely for species with low abundances (e.g., see points on the bottom-right of both panels of Fig. 5c). These deviations stem from several sources of error: primarily measurement errors and biological variability, alongside uncertainty in fitted interaction strengths, as well as GLV dynamics being an approximation of the true dynamics. By inferring invasion fitness using GLV models fitted for these datasets (Methods), we also confirmed that dressed invasion fitness is a much better predictor of invader abundance than naive invasion fitness (Fig. S9 and Fig. S10). Taken together, these results demonstrate that our framework can not only predict, but also provide insight into invasion outcomes for diverse ecosystems spanning a range of environments, size scales, and trophic levels.
@@ VI Discussion
A fundamental insight from our theory is the identification of a “dressed invasion fitness” as the central quantity governing invasion outcomes. Unlike conventional “naive” invasion fitness [34, 47] which only measures the initial growth rate of an invader in the resident community, the dressed invasion fitness incorporates ecological feedbacks between the invader and the resident community. We show that this distinction is crucial for accurately predicting both invasion success and the resulting community structure. This is because mathematically, the effect of an invader abundance decomposes into a contribution proportional to its dressed invasion fitness and a contribution from invasion-induced extinctions. Both these contributions incorporate ecological feedbacks between the invader and the full community into which it invades. Similar feedbacks appear in the cavity method for models of ecosystems, but these approaches are completely different in scope. Our framework uses known interactions to make predictions for specific invasions in specific communities, while the cavity method uses random parameters and can only describe statistical properties of an ensemble of invasions.
A direct implication of our framework is that eco-evolutionary dynamics in complex communities follow a pattern where successful mutants predominantly replace low-abundance community members rather than their parent species [10]. This stands in contrast to the canonical view from population genetics — where displacement occurs among closely related lineages — as well as pairwise niche overlap metrics [14]. While parents and mutants do compete more strongly among themselves than with other community members, the cumulative effect of weaker interactions across a sufficiently diverse community may ultimately also affect the fate of a mutant. In our work, this occurs because mutants and parents interact not only directly with each other, but also through global feedbacks mediated by the entire community structure. This contrasting perspective motivates further empirical tests as well as theoretical work on population genetics in complex ecological communities. After successful invasion, the extinction probabilities of other residents decay exponentially with their pre-invasion abundance, effectively shielding abundant species while rendering rare species vulnerable regardless of their relatedness to the invader. These patterns are consistent with observations from natural communities which show a decreasing pattern of coexistence with phenotypic similarity [43, 48]. Our results augment these observations: despite the decrease, in a diverse community, phenotypically similar species are still reasonably likely ($\sim 40\%$) to coexist (Fig. S11). This observation also highlights how ecological feedbacks fundamentally reshape evolutionary trajectories in ways that cannot be captured by traditional models that ignore invader-induced feedback.
@@ Acknowledgments
Appendix A Derivation of linear response theory for the Generalized Lotka-Volterra (GLV) models
@@ A.1 Primer: Environmental perturbations without extinctions
$\vec{X}^{\text{new}}_{\mathcal{S}}$ $=A_{\mathcal{S}\mathcal{S}}^{-1}(\vec{r_{\mathcal{S}}}-A_{\mathcal{S}0}{X}_{0}).$ (17)
@@
$\vec{r_{\mathcal{S}}}-A_{\mathcal{S}\mathcal{S}}\vec{X}^{\text{old}}_{\mathcal{S}}-A_{\mathcal{S}\mathcal{E}}\vec{X}_{\mathcal{E}}$ $=0,$ (18)
$\vec{X}^{\text{old}}_{\mathcal{S}}$ $=A_{\mathcal{S}\mathcal{S}}^{-1}(\vec{r_{\mathcal{S}}}-A_{\mathcal{S}\mathcal{E}}\vec{X}_{\mathcal{E}}).$ (19)
@@
$\delta\vec{X}_{\mathcal{S}}$ $=\vec{X}^{\text{new}}_{\mathcal{S}}-\vec{X}^{\text{old}}_{\mathcal{S}},$ (20)
$=A_{\mathcal{S}\mathcal{S}}^{-1}(-A_{\mathcal{S}0}{X}_{0}+A_{\mathcal{S}\mathcal{E}}\vec{X}_{\mathcal{E}}).$ (21)
@@
Figure S1: Equations showing how our theory partitions an ecosystem into invaders, survivors and extinct species. (a) shows how the structure of interactions changes in the Lotka-Volterra model before and after an invasion, and (b) shows how the community shifts and invader abundance is related to each part of the system shown in Eq. (26)
@@ A.4 Extended framework for multiple invaders and environmental perturbations
$\vec{r}^{\text{inv}}=-A_{\mathcal{S}\mathcal{I}}\vec{X}_{\mathcal{I}},\quad\vec{r}^{\text{ext}}=A_{\mathcal{S}\mathcal{E}}\vec{X}_{\mathcal{E}}$ (28)
@@
Substituting (29) into (30) like before, we obtain the generalized version of Eq. (5)
@@
Appendix B Predicting invasion in the presence of multi-stability
Our framework can predict the correct invasion outcome even if the dynamical system has multiple stable fixed points, given that identity of surviving species are not significantly changed. We demonstrate this using Lotka-Volterra model with high enough variance in interaction strength. It has been shown that a phase transition change the system from unique stable steady states to multiple stable steady states [54] in random Lotka-Volterra models when the variance of insteraction increases. We demonstrate this by explicitly sampling many initial state to check if the system exhibit multiple fixed points, then estimates the portion of steady-states exhibiting multistability given the level of variance. Finally we choose a high enough variance that guarantee multi-stability happens more than 80 percent of the time.
In the main text, we sample the invader using the same distribution as that of the regional species pool, which forms the resident community before invasion. This choice models the natural occurrence of invaders originating from a statistically similar pool of species. However, when we increase the variance of interactions in the regional species pool, some interactions become negative, representing “cooperative” rather than competitive interactions. This cooperation allows species from the pool to become large-impact invaders that can potentially drive most resident species to extinction. While the self-consistency equation remains exact in such situations, the iterative scheme assumes that a considerable fraction of species survive after invasion. Therefore, in the demonstration in Fig. S2, we reduced the variance of the interaction coefficients for the pool from which we sample invaders. In this simulation, we set the mean interaction to be $\mu_{A}=0.1$, the variance of the regional species pool for the resident community as $\sigma_{A}=3.16$, and the variance for sampling invaders as $\sigma_{A}^{\text{invader}}=0.316$. The size of the regional species pool is 10, with approximately 3 species surviving in the resident community. This demonstrates that the self-consistency equations and the iterative approach remain applicable in the presence of multiple steady states.
Figure S2: Comparison of simulation outcomes and theoretical predictions for Lotka-Volterra models exhibiting multiple stable equilibria prior to invasion. The resident community’s interaction matrix is sampled with high variance ($\sigma_{c}=3.16$), generating systems with multiple locally stable fixed points identified by simulating dynamics from diverse initial conditions. Invader–community interaction strengths are drawn from a distribution with lower variance ($\sigma_{\text{inv}}=0.316$), ensuring that invaders exert a limited perturbative effect on the resident community.
Appendix C General linear response theory
@@
where $X_{i}$ represents the abundance of each species, and $\vec{r}$ parameterizes the environmental influences on growth. The function $f_{i}(X_{i})$ governs species extinction dynamics. It is positive for surviving species, and it captures the concept of extinction for biotic species as $f_{i}(X_{i})$ must satisfy
$f_{i}(0)=0.$ (35)
@@
where the second condition corresponds to the stability to re-introduction by extinct species. This tells us that instead of linearizing the complete dynamics in Eq.(34), it is more natural to linearize the steady-state condition if we consider perturbations to the ecosystem. Note that the full dynamics are always nonlinear because of the presence of $f_{i}(X_{i})$. When we refer to linearization in this appendix, we refer to approximating $g_{i}$ as a linear function by expanding it around a given steady-state set of abundances $\vec{X}$ and growth rates $\vec{r}$. As we show below, this approximation is mathematically equivalent to assuming that the effective interaction matrix strengths do not change too much between the pre- and post-invasion communities. We might refer to this general class of models with nonlinear $g_{i}$ as “nonlinear ecological models” for brevity; to clarify all models considered in this manuscript are nonlinear in their full dynamics. We merely mean that the per-capita growth rates are nonlinear when we refer to this shorthand.
We first study the scenario where all species survive, following the first condition in Eq.(36). We want to study the abundance change to each species, $\vec{X}\to\vec{X}+\delta\vec{X}$, in response to perturbation of environmental parameter $\vec{r}\to\vec{r}+\delta\vec{r}$. The new steady state needs to satisfy
@@
Define the effective interaction matrices as the derivatives of the non-zero steady state conditions,
$\mathcal{A}_{ij}$ $=\left.\frac{\partial g_{i}(\vec{X},\vec{r})}{\partial X_{j}}\right|_{\vec{X}=\vec{X}^{\text{old}}},\quad\mathcal{B}_{ij}=\left.\frac{\partial g_{i}(\vec{X},\vec{r})}{\partial r_{j}}\right|_{\vec{X}=\vec{X}^{\text{old}}}$ (38)
Note that in general effective interactions $\mathcal{A}$ and $\mathcal{B}$ depend on $\vec{X}$, but assuming they remain constant at the new steady state $\vec{X}^{*}$ is precisely the same as performing a first-order Taylor expansion of the new steady state condition in Eq. (37). This linear approximation of the change in the steady state function around $\vec{X}^{*}$ yields
$\mathcal{A}\delta\vec{X}+\mathcal{B}\delta\vec{r}=0.$ (39)
By simplifying and labeling
$A^{-1}=-\mathcal{A}^{-1}\mathcal{B},$
the abundance shift satisfies the linear approximation
$\delta\vec{X}=A^{-1}\delta\vec{r}.$ (40)
Here, the linearity expressed through the derivatives $\mathcal{A}$ and $\mathcal{B}$ corresponds exactly to the statement that changes in $\vec{X}$ can be well approximated by applying these effective interaction matrices to changes in $\vec{r}$, illustrating the equivalence between locally linear functions and their differential approximations.
In the context of invasion theory, we have effective environmental perturbation $\delta\vec{r}^{\text{eff}}$ consists of invasion, extinction and actual environmental perturbation, the same way as linear model in Eq.(26) and Fig. S1:
where
$\delta\vec{r}^{\text{inv}}=-{A}_{\mathcal{S}\mathcal{I}}\vec{X}_{\mathcal{I}},\quad\delta\vec{r}^{\text{ext}}={A}_{\mathcal{S}\mathcal{E}}\vec{X}_{\mathcal{E}}.$ (42)
Therefore the abundance shift to surviving species is
$\delta X={A}^{-1}_{\mathcal{S}\mathcal{S}}(-{A}_{\mathcal{S}\mathcal{I}}\vec{X}_{\mathcal{I}}+{A}_{\mathcal{S}\mathcal{E}}\vec{X}_{\mathcal{E}}+\delta\vec{r})$ (43)
To summarize, to predict the linear response of the steady state in any nonlinear model, the first step is to compute the effective interaction matrix, and compute the matrix inverse given by $A^{-1}$. Once this matrix is determined, the problem simplifies and becomes equivalent to the generalized Lotka-Volterra model. This simple exercise of Taylor expansion justifies the use of the Lotka-Volterra model despite its simplicity - it captures the linear response of ecosystems near their steady states. Finally, the result maps to the one derived for the generalized Lotka-Volterra model in Eq.(6).
A similar derivation for Eq.(5) will require solving only the equation for invader steady state abundance, i.e., $g_{i\in\mathcal{I}}(\vec{X}^{\text{new}},r_{i})=0$.
$g_{i\in\mathcal{I}}(\vec{X}_{\mathcal{S}}^{\text{old}}+\delta\vec{X},\vec{X}_{\mathcal{E}},X_{i},r_{i})$ $=0,$ (44)
$\sum_{j\in\mathcal{S}}\frac{\partial g_{i\in\mathcal{I},j}}{\partial X_{j}}\delta{x_{j}}+\sum_{k\in\mathcal{I}}\frac{\partial g_{i\in\mathcal{I},k}}{\partial X_{k}}{x_{k}}+g_{i\in\mathcal{I}}(\vec{X}^{\text{old}}_{\mathcal{S}},\vec{X}^{\text{old}}_{\mathcal{E}}=0,X_{i}=0,r_{i})$ $=0.$ (45)
The last part of the expression corresponds almost exactly to the naive invasion fitness, except that contributions from extinct species are set to zero. This portion uniquely captures the nonlinear effects present in the growth rate, whereas the first two terms arise purely from the linear response of the system.
In matrix notation, it becomes simply
${A}_{IS}\delta{\vec{X}_{S}}+{A}_{\mathcal{I}\mathcal{I}}\vec{X}_{\mathcal{I}}+\vec{g}_{\mathcal{I}}(\vec{X}^{\text{old}},X_{i}=0,r_{0})$ $=0.$ (46)
Therefore, the invader abundance formula in terms of dressed invasion fitness Eq.(10) applies even for nonlinear consumer growth rate.
Appendix D Extinct species have negative invasion fitness
@@
$\frac{d}{dt}(x_{i}+\delta x_{i})=(x_{i}+\delta x_{i})(k_{i}-\sum_{j\neq i}A_{ij}x_{j}-A_{ii}\delta x_{i})$ (47)
@@
$\frac{d\delta x_{i}}{dt}=\delta x_{i}(k_{i}-\sum_{j\neq i}A_{ij}x_{j}-A_{ii}\delta x_{i})$ (48)
@@
$\frac{d\delta X_{i}}{dt}=g_{i}(\vec{X},\vec{r})\frac{\partial f_{i}}{\partial X_{i}}\delta X_{i}+f_{i}(X_{i})\frac{\partial g_{i}}{\partial X_{i}}\delta X_{i}$ (50)
@@
$\left.g_{i}(\vec{X},\vec{r})\frac{\partial f_{i}}{\partial X_{i}}\right|_{X_{i}=0}<0+\left.f_{i}(X_{i})\frac{\partial g_{i}}{\partial X_{i}}\right|_{X_{i}=0}<0$ (51)
@@
Appendix E Derivation of linear response in CRMs